OpenStudy (anonymous):

Quadratic formula: show work please!!!! 9x^2+48x+64

jimthompson5910 (jim_thompson5910):

\[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(48)\pm\sqrt{(48)^2-4(9)(64)}}{2(9)}\] \[\Large x = \frac{-48\pm\sqrt{2304-(2304)}}{18}\] \[\Large x = \frac{-48\pm\sqrt{0}}{18}\] \[\Large x = \frac{-48+\sqrt{0}}{18} \ \text{or} \ x = \frac{-48-\sqrt{0}}{18}\] \[\Large x = \frac{-48+0}{18} \ \text{or} \ x = \frac{-48-0}{18}\] \[\Large x = \frac{-48}{18} \ \text{or} \ x = \frac{-48}{18}\] \[\Large x = \frac{-48}{18}\] \[\Large x = -\frac{8}{3}\] So the only solution is \[\Large x = -\frac{8}{3}\]

OpenStudy (anonymous):

\[-b \pm \sqrt{b ^{2}-4ac}/2a\]

OpenStudy (anonymous):

a=9; b=48; c=64... plug in & solve